There will be an access day (2.5-4 hours) tomorrow Wednesday. There are a number of issues to be fixed for Brahms learned during the store today. Due to the limited time some may have to wait for the May 31 access day. . HV 4032 #3-6+16 cannot communicate, making most of the exp inoperable. . one L 1440 in FFS does not communicate apart from the first few commands. . The third section of T4 seems not to work (see e-log). . A lecroy TDC in FB3 has failed (slat 12) should be changed. . the v124 bunchcounter module has problems. . some CC channels seems not working - should be chekced in IR. the basic trigger timing seems ok. Extensive checking could not be done due to the lack of HV on multiple systems. Comments and helpful advice is appreciated. regards Flemming . -------------------------------------------- Flemming Videbaek Physics Department Bldg 510-D Brookhaven National Laboratory Upton, NY11973 phone: 631-344-4106 cell: 631-681-1596 fax: 631-344-1334 e-mail: videbaek @ bnl.gov _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Tue May 23 19:22:45 2006
This archive was generated by hypermail 2.1.8 : Tue May 23 2006 - 19:22:59 EDT